a dynamo delivers 30.0 a at 120 v to an external circuit when operating at 1200 rpm. what torque is required to drive the generator at this speed if the total power losses are 400 W?e

1
Expert's answer
2021-11-24T19:14:00-0500

I = 30.0 A

V = 120 V

"P = IV = 30.0 \times 120 = 3600 \;W"

Power loss = 400 W

Total power output = 3600 + 400 = 4000 W = 4 kW